OpenStudy (anonymous):

find all points on the curve y=x^2-2x+4 for which the tangent line passes through the point (1,-1)

OpenStudy (amistre64):

detemine the slope of the line at some point (a , b) y' = 2x - 2, the slope at x=a is simply 2a-2 y* = (2a-2)(x-a) + (b) y* = (2a-2)(x-a) + (a^2-2a+4) will have to work for the point (1,-1) -1 = (2a-2)(1-a) + (a^2-2a+4) , solve for a